# Ramification filtration via deformations, II

### v3

@@ 2.4. Lie algebras $\mathcal{N}^{{\dagger}}$ and $\bar{\mathcal{N}}^{{\dagger}}$
For $s\geqslant 0$, ${\mathcal{N}}^{\,{\dagger}}(s):=\oplus_{j+m\geqslant s}{\mathcal{N}}^{\,{\dagger}}[j,m]$ form a decreasing filtration in ${\mathcal{N}}^{\,{\dagger}}$ such that $[{\mathcal{N}}^{{\dagger}}(s_{1}),{\mathcal{N}}^{{\dagger}}(s_{2})]\subset{\mathcal{N}}^{\,{\dagger}}(s_{1}+s_{2})$.
Similarly to Section 1.5 we can introduce for $w\geqslant 1$, a closed submodule $\mathcal{L}^{{\dagger}}_{\mathcal{N}}(w)$ in $\mathcal{L}^{{\dagger}}$ generated by all $p^{m}[D_{\bar{\iota}\bar{n}}]$ with $|\bar{\iota}|\geqslant p^{M-1}b^{*}U(w-1,m)$. Then there are analogs of Proposition 1.6 and Corollary 1.8
$\mathcal{L}^{{\dagger}}_{\mathcal{N}}(w)\cap C_{2}(\mathcal{L}^{{\dagger}}_{W_{M}(k)})\subset\sum_{w_{1}+w_{2}\geqslant w}[\mathcal{L}_{\mathcal{N}}(w_{1}),\mathcal{L}_{\mathcal{N}}(w_{2})],\ \ \mathcal{L}^{{\dagger}}_{\mathcal{N}}(w)\subset\mathcal{L}^{{\dagger}}(w)\,.$
As a result, for any $w$ and $\mathrm{m}^{\prime}=W_{M}(k)[[t_{1}]]$, it holds
$\mathcal{N}^{{\dagger}}(s)\subset\mathcal{L}^{{\dagger}}_{\mathrm{m}^{\prime}}+\mathcal{L}^{{\dagger}}(1+s)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Remark.
Similar definitions can be done in the context of the Lie algebra $\bar{\mathcal{L}}^{{\dagger}}$. In particular, we obtain an analog $\bar{\mathcal{N}}^{{\dagger}}$ of $\mathcal{N}^{{\dagger}}$ with corresponding filtration $\bar{\mathcal{N}}^{{\dagger}}(s)$ and grading $\bar{\mathcal{N}}^{{\dagger}}[j,m]$. The projection $\mathcal{N}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}$ allows us to transform above properties to $\bar{\mathcal{N}}^{{\dagger}}$.
In particular, we obtain
Corollary 2.5.
$\bar{\mathcal{N}}^{\,{\dagger}}(p-1)\subset\mathcal{L}^{\,{\dagger}}_{\mathrm{m}^{\prime}}$.
Definition.
Set $\widetilde{\mathcal{N}}^{\,{\dagger}}=\bar{\mathcal{N}}^{\,{\dagger}}/\bar{\mathcal{N}}^{\,{\dagger}}(p)$.
The above grading on $\bar{\mathcal{N}}^{\,{\dagger}}$ induces the grading $\widetilde{\mathcal{N}}^{\,{\dagger}}=\oplus_{j,m}{\widetilde{\mathcal{N}}}^{\,{\dagger}}[j,m]$. Note that $\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]=0$ if $j+m\geqslant p$.
@@ 2.5. The action $\Omega$
For $b=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$ and $l=p^{A}t_{1}^{-b}D^{\,{\dagger}}_{\bar{\iota}\bar{n}}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\,$, set
@@
where $E(X)$ is the Artin-Hasse exponential.
Note that $E(t_{1}^{b^{*}})^{(p^{m}\alpha)p^{M-1}}\equiv\widetilde{\exp}\left((p^{m}\alpha)S_{b^{*}}\right)\mathrm{mod}\,S_{b^{*}}^{p}$ where $S_{b^{*}}=t_{1}^{b^{*}p^{M-1}}+pt_{1}^{b^{*}p^{M-2}}+\ldots+p^{M-1}t_{1}^{b^{*}}$ and $\widetilde{\exp}$ is the truncated exponential. Therefore, in $\widetilde{\mathcal{N}}^{\,{\dagger}}$ it holds $\Omega(l)=\widetilde{\exp}\left((p^{m}\alpha)S_{b^{*}}\right)l$.
@@
Note that $S_{b^{*}}^{p}\bar{\mathcal{N}}^{{\dagger}}\subset\bar{\mathcal{N}}^{{\dagger}}(p)$. Therefore, it is essential that we could define the action $\Omega$ only in the context of $\widetilde{\mathcal{N}}^{{\dagger}}=\bar{\mathcal{N}}^{{\dagger}}/\bar{\mathcal{N}}^{{\dagger}}(p)$. However, $\widetilde{\mathcal{N}}^{{\dagger}}$ is still related to the Lie algebra $\bar{\mathcal{L}}^{{\dagger}}$, i.e. the natural embedding ¯$\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}\subset\bar{\mathcal{L}}^{{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ still induces embedding ¯$\bar{\mathcal{L}}^{{\dagger}}\to\widetilde{\mathcal{N}}^{{\dagger}}$ after taking the projection from $\bar{\mathcal{N}}^{{\dagger}}$ to $\widetilde{\mathcal{N}}^{{\dagger}}$. (Use that $\bar{\mathcal{N}}^{{\dagger}}(p)\subset\bar{\mathcal{L}}^{{\dagger}}_{\mathrm{m}^{\prime}}$.)
By linearity, $\Omega$ can be extended to the whole $\widetilde{\mathcal{N}}^{\,{\dagger}}$.
The map $n\mapsto\Omega_{\gamma}(n):=\Omega^{\gamma}(n)$, $\gamma\in\mathbb{Z}/p^{M}$, is a well-defined action of the additive group $\mathbb{Z}/p^{M}$ on the Lie algebra $\widetilde{\mathcal{N}}^{\,{\dagger}}$. This action is unipotent: for any $n\in\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]$, $\Omega_{\gamma}(n)-n\in\oplus_{i>j}\widetilde{\mathcal{N}}^{\,{\dagger}}[i,m]\,.$
@@
(2.1) $\ \ \bar{e}^{\,{\dagger}}\equiv\sum\limits_{\iota\in\mathfrak{A}^{0}(p)}t_{1}^{-\iota}D^{\,{\dagger}}_{\iota 0}\,\mathrm{mod}\,C_{2}(\bar{\mathcal{L}}^{{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})})\,.$
@@
Such element is used usually in the nilpotent Artin-Shreier theory, cf. Section 3. In the above expression the elemnt $D^{{\dagger}}_{00}$ is not defined. In this and similar case for the element $e$ from Section 3 we set $D^{{\dagger}}_{00}:=\alpha D^{{\dagger}}_{0}$, where $\alpha=\hat{\alpha}\,\mathrm{mod}\,p^{M}\in W_{M}(k)$ and the trace of $\hat{\alpha}$ over $\mathbb{Q}_{p}$ is equal to 1.
@@
A choice of $\bar{e}^{\,{\dagger}}$ allows us to associate to the above defined action $\Omega_{\gamma}$ the “conjugated” action of $A^{\,{\dagger}}_{\gamma}$ on $\bar{\mathcal{L}}^{\,{\dagger}}$ as follows.
Let $\widetilde{\mathcal{N}}^{\,{\dagger}}\langle j\rangle=\oplus_{m}\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]$ and $\widetilde{\mathcal{N}}^{\,{\dagger}}\{s\}:=\oplus_{j\geqslant s}\widetilde{\mathcal{N}}^{\,{\dagger}}\langle j\rangle$.
@@
For $\gamma\in\mathbb{Z}/p^{M}$, there are unique $\widetilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}$ and $A^{\,{\dagger}}_{\gamma}\in\mathrm{Aut}_{\mathrm{Lie}}\bar{\mathcal{L}}^{\,{\dagger}}$ such that
a) $\sigma\tilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}$ and $\Omega_{\gamma}(\widetilde{e}^{\,{\dagger}})=(\sigma\widetilde{c}_{\gamma})\circ(A^{\,{\dagger}}_{\gamma}\otimes\mathrm{id}_{\mathcal{K}})\widetilde{e}^{\,{\dagger}}\circ(-\widetilde{c}_{\gamma})$;
b) for any $\iota\in\mathfrak{A}^{0}(p)$, $A^{\,{\dagger}}_{\gamma}(D^{\,{\dagger}}_{\iota 0})-D^{\,{\dagger}}_{\iota 0}\in\oplus_{m<\mathrm{ch}(\iota)}\bar{\mathcal{L}}^{\,{\dagger}}[m]_{W_{M}(k)}$.
@@
Suppose $n\in\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ where $j+m\geqslant 1$. Then there are unique $\mathcal{S}(n),\mathcal{R}(n)\in\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ such that:
a) $\mathcal{R}(n)=\sum\limits_{\iota}t^{-\iota}l_{\iota}$ where $\iota$ runs over $\mathfrak{A}^{+}(p)$ with $\mathrm{ch}(\iota)=j+m$ and all $l_{\iota}\in\left(\bar{\mathcal{L}}^{\,{\dagger}}[m]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})\right)_{W_{M}(k)}$;
@@
It will be sufficient to specify the elements $\mathcal{R}(n)$ and $\mathcal{S}(n)$ for $n=p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where $b\in\mathfrak{A}^{0}$, $\mathrm{ch}(b)=j+m$, $\mathrm{ch}(\bar{\iota})=m$.
@@
If $b>0$ we set $\mathcal{R}(n)=0$ and $\mathcal{S}(n)=-\sum_{i\geqslant 0}\sigma^{i}(n)$.
If $b=\iota_{0}p^{i_{0}}<0$ with $\iota_{0}\in\mathfrak{A}^{+}(p)$ we set $\mathcal{R}(n)=\sigma^{-i_{0}}n$ and $\mathcal{S}(n)=\sum_{0\leqslant j<i_{0}}\sigma^{-j}n$.
The uniqueness property holds because $\mathrm{ch}(b)\geqslant 1$ implies that $b\neq 0$. Lemma is proved. ∎
Continue the proof of Prop. 2.6.
Use induction on $i\geqslant 1$ to prove the proposition modulo the ideals $\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap C_{i}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$.
— If $i=1$ take $\widetilde{c}_{\gamma}=0$, $A^{\,{\dagger}}_{\gamma}=\mathrm{id}$ and use that $\Omega_{\gamma}(\widetilde{e}^{\,{\dagger}})-\widetilde{e}^{\,{\dagger}}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\,.$
— Suppose $i\geqslant 2$, $\widetilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{{\dagger}}\{1\}$ and $A_{\gamma}^{\,{\dagger}}\in\mathrm{Aut}_{\mathrm{Lie}}(\bar{\mathcal{L}}^{\,{\dagger}})$ are such that
$H=\Omega_{\gamma}\widetilde{e}^{\,{\dagger}}-(\sigma\widetilde{c}_{\gamma})\circ(A^{\,{\dagger}}_{\gamma}\otimes\mathrm{id}_{\mathcal{K}^{\prime}})\widetilde{\,e}^{\,{\dagger}}\circ(-\widetilde{c}_{\gamma})\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap\,C_{i}(\widetilde{\mathcal{N}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Then $\mathcal{R}(H),\mathcal{S}(H)\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap\,C_{i}(\widetilde{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime}}$.
Set $\mathcal{R}(H)=\sum\limits_{\mathrm{ch}(\iota)\geqslant 1+m}t^{-\iota}H_{\iota m}$, where all $H_{\iota m}\in\widetilde{\mathcal{L}}^{\,{\dagger}}[m]_{k}\cap C_{i}(\bar{\mathcal{L}}^{\,{\dagger}})_{W_{M}(k)}$.
@@
$\Omega_{\gamma}\widetilde{e}^{\,{\dagger}}\equiv(\sigma\widetilde{c}^{\,\prime}_{\gamma 1})\circ(A^{\,{\dagger}\prime}_{\gamma}\ \otimes\mathrm{id}_{\mathcal{K}})\widetilde{e}^{\,{\dagger}}\circ(-\widetilde{c}^{\,\prime}_{\gamma})\,\mathrm{mod}\,\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap C_{i+1}(\widetilde{\mathcal{N}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
@@
We have obviously the following properties.

### v5

@@ 2.3. Lie algebras ${\mathcal{L}}^{\,{\dagger}}$ and $\bar{\mathcal{L}}^{\,{\dagger}}$
Let $\widetilde{\mathcal{N}}^{\,{\dagger}}$ be the submodule of $\bar{\mathcal{N}}^{\,{\dagger}}/\bar{\mathcal{N}}^{\,{\dagger}}\langle p-1\rangle$ such that its $W(k)$-extension of scalars is generated by the elements of the form $t^{-\jmath}[D^{{\dagger}}_{\bar{\iota}\bar{n}}]$, where $\jmath\in\mathfrak{A}^{0}$.
Clearly, $\widetilde{\mathcal{N}}^{\,{\dagger}}$ is a Lie subalgebra of $\bar{\mathcal{N}}^{\,{\dagger}}/\bar{\mathcal{N}}^{\,{\dagger}}(p-1)$ and it has a double grading via submodules $\widetilde{\mathcal{N}}^{\,{\dagger}}[m,t]$ generated by the elements of the form $t^{-j}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where $\mathrm{ch}(j)=t$ and $\mathrm{ch}(\bar{\iota})=m$. Here $20\leqslant m\leqslant t\leqslant p-$ 2. This Lie algebra plays a key role in our further interpretation of the ramification criterion.
@@ 2.5. The action $\Omega$
For $b=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$ and $l=p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]\in\widetilde{\mathcal{N}}^{\,{\dagger}}\,$, set
@@
where $E(X)$ is the Artin-Hasse exponential. We have
$E(t_{1}^{b^{*}})^{(p^{m}\alpha)p^{M-1}}\equiv\widetilde{\exp}\left((p^{m}\alpha)S_{b^{*}}\right)\mathrm{mod}\,S_{b^{*}}^{p}\,,$
where $S_{b^{*}}=t_{1}^{b^{*}p^{M-1}}+pt_{1}^{b^{*}p^{M-2}}+\ldots+p^{M-1}t_{1}^{b^{*}}$ and $\widetilde{\exp}$ is the truncated exponential. Therefore, $\Omega(l)=\widetilde{\exp}\left((p^{m}\alpha)S_{b^{*}}\right)l$.
@@
Note that $S_{b^{*}}^{p}\bar{\mathcal{N}}^{{\dagger}}\subset\bar{\mathcal{N}}^{{\dagger}}(p)$. So, it is essential that we could define the action $\Omega$ in the context of $\widetilde{\mathcal{N}}^{{\dagger}}$ modulo $\bar{\mathcal{N}}^{{\dagger}}(p-1)$ (which contains $\bar{\mathcal{N}}^{\,{\dagger}}(p)$). At the same time, $\widetilde{\mathcal{N}}^{{\dagger}}$ is still sufficient to recover the Lie algebra $\bar{\mathcal{L}}^{{\dagger}}$. Indeed, the natural embedding ¯$\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}\subset\bar{\mathcal{L}}^{{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ still induces embedding ¯$\bar{\mathcal{L}}^{{\dagger}}\to\widetilde{\mathcal{N}}^{{\dagger}}$ after taking the projection from $\bar{\mathcal{N}}^{{\dagger}}$ to $\widetilde{\mathcal{N}}^{{\dagger}}$. (Use that $\bar{\mathcal{N}}^{{\dagger}}(p-1)\subset\bar{\mathcal{L}}^{{\dagger}}_{\mathrm{m}^{\prime}}$.)
By linearity, $\Omega$ (when considered in the context of $\widetilde{\mathcal{N}}^{{\dagger}}$) can be extended to the whole $\widetilde{\mathcal{N}}^{\,{\dagger}}$.
The map $n\mapsto\Omega_{\gamma}(n):=\Omega^{\gamma}(n)$, $\gamma\in\mathbb{Z}/p^{M}$, is a well-defined action of the additive group $\mathbb{Z}/p^{M}$ on the Lie algebra $\widetilde{\mathcal{N}}^{\,{\dagger}}$. This action is unipotent: for any $n\in\widetilde{\mathcal{N}}^{\,{\dagger}}[m,t]$, $\Omega_{\gamma}(n)-n\in\oplus_{t^{\prime}>t}\widetilde{\mathcal{N}}^{\,{\dagger}}[m,t^{\prime}]\,.$
@@
(2.2) $\ \ \bar{e}^{\,{\dagger}}\equiv\sum\limits_{\iota\in\mathfrak{A}^{0}(p)}t_{1}^{-\iota}D^{\,{\dagger}}_{\iota 0}\,\mathrm{mod}\,C_{2}(\bar{\mathcal{L}}^{{\dagger}}_{O_{M}^{\prime}(\mathcal{K}^{\prime})})\,.$
@@
Such an element is used usually in the nilpotent Artin-Shreier theory, cf. Section 3. In the above expression the element $D^{{\dagger}}_{00}$ is not defined. In this and similar cases for the element $e$ from Section 3 we set $D^{{\dagger}}_{00}:=\alpha D^{{\dagger}}_{0}$, where $\alpha=\hat{\alpha}\,\mathrm{mod}\,p^{M}\in W_{M}(k)$ and the trace of $\hat{\alpha}$ over $\mathbb{Q}_{p}$ is equal to 1.
@@
A special choice of $\bar{e}^{\,{\dagger}}$ allows us to associate to the above defined action $\Omega_{\gamma}$ on $\widetilde{\mathcal{N}}^{{\dagger}}$ the “conjugated” action $A^{\,{\dagger}}_{\gamma}$ on $\bar{\mathcal{L}}^{\,{\dagger}}$ as follows.
@@
For $\gamma\in\mathbb{Z}/p^{M}$, there are unique $\widetilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\langle 1\rangle$ and $A^{\,{\dagger}}_{\gamma}\in\mathrm{Aut}_{\mathrm{Lie}}\bar{\mathcal{L}}^{\,{\dagger}}$ such that
a) $\sigma\tilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\langle 1\rangle$ and $\Omega_{\gamma}(\widetilde{e}^{\,{\dagger}})=(\sigma\widetilde{c}_{\gamma})\circ(A^{\,{\dagger}}_{\gamma}\otimes\mathrm{id}_{\mathcal{K}})\widetilde{e}^{\,{\dagger}}\circ(-\widetilde{c}_{\gamma})$;
b) for any $\iota\in\mathfrak{A}^{0}(p)$, $A^{\,{\dagger}}_{\gamma}(D^{\,{\dagger}}_{\iota 0})-D^{\,{\dagger}}_{\iota 0}\in\underset{m<\mathrm{ch}(\iota)}{\oplus}\bar{\mathcal{L}}^{\,{\dagger}}[m]_{W(k)}$.
@@
Suppose $n\in\widetilde{\mathcal{N}}^{\,{\dagger}}[m,t]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}(\mathcal{K}^{\prime})}$ where $t\geqslant 1$. Then there are unique $\mathcal{S}(n),\mathcal{R}(n)\in\widetilde{\mathcal{N}}^{\,{\dagger}}[m,t]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}(\mathcal{K}^{\prime})}$ such that:
a) $\mathcal{R}(n)=\sum\limits_{\iota}t^{-\iota}l_{\iota}$ where $\iota$ runs over $\mathfrak{A}^{+}(p)$ with $\mathrm{ch}(\iota)=t$ and all $l_{\iota}\in\left(\bar{\mathcal{L}}^{\,{\dagger}}[m]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})\right)_{W(k)}$;
@@
It will be sufficient to specify the elements $\mathcal{R}(n)$ and $\mathcal{S}(n)$ for $n=p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where $b\in\mathfrak{A}^{0}$, $\mathrm{ch}(b)=t$ and $\mathrm{ch}(\bar{\iota})=m$.
@@
If $-b>0$ we set $\mathcal{R}(n)=0$ and $\mathcal{S}(n)=-\sum_{i\geqslant 0}\sigma^{i}(n)$.
If $-b=-\iota_{0}p^{i_{0}}<0$ with $\iota_{0}\in\mathfrak{A}^{+}(p)$ we set $\mathcal{R}(n)=\sigma^{-i_{0}}n$ and $\mathcal{S}(n)=\sum_{0\leqslant j<i_{0}}\sigma^{-j}n$.
The uniqueness property holds because $\mathrm{ch}(b)\geqslant 1$ implies that $b\neq 0$. The lemma is proved. ∎
Continue the proof of Prop. 2.8.
Use induction on $i\geqslant 1$ to proceed modulo the ideals $\widetilde{\mathcal{N}}^{\,{\dagger}}\langle 1\rangle\cap C_{i}(\bar{\mathcal{L}}^{\,{\dagger}})_{O_{M}^{\prime}(\mathcal{K}^{\prime})}$.
— If $i=1$ take $\widetilde{c}_{\gamma}=0$, $A^{\,{\dagger}}_{\gamma}=\mathrm{id}$ and use that $\Omega_{\gamma}(\widetilde{e}^{\,{\dagger}})-\widetilde{e}^{\,{\dagger}}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\langle 1\rangle\,.$
— Suppose $i\geqslant 2$, $\widetilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{{\dagger}}\langle 1\rangle$ and $A_{\gamma}^{\,{\dagger}}\in\mathrm{Aut}_{\mathrm{Lie}}(\bar{\mathcal{L}}^{\,{\dagger}})$ are such that
$H=\Omega_{\gamma}\widetilde{e}^{\,{\dagger}}-(\sigma\widetilde{c}_{\gamma})\circ(A^{\,{\dagger}}_{\gamma}\otimes\mathrm{id}_{\mathcal{K}^{\prime}})\widetilde{\,e}^{\,{\dagger}}\circ(-\widetilde{c}_{\gamma})\in\widetilde{\mathcal{N}}^{\,{\dagger}}\langle 1\rangle\cap\,C_{i}(\widetilde{\mathcal{N}}^{\,{\dagger}})_{O_{M}^{\prime}(\mathcal{K}^{\prime})}\,.$
Then $\mathcal{R}(H),\mathcal{S}(H)\in\widetilde{\mathcal{N}}^{\,{\dagger}}\langle 1\rangle\cap\,C_{i}(\widetilde{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})})$.
Set $\mathcal{R}(H)=\sum\limits_{\mathrm{ch}(\iota)\geqslant 1+m}t^{-\iota}H_{\iota m}$, where all $H_{\iota m}\in\widetilde{\mathcal{L}}^{\,{\dagger}}[m]_{k}\cap C_{i}(\bar{\mathcal{L}}^{\,{\dagger}})_{W(k)}$.
@@
$\Omega_{\gamma}\widetilde{e}^{\,{\dagger}}\equiv(\sigma\widetilde{c}^{\,\prime}_{\gamma 1})\circ(A^{\,{\dagger}\prime}_{\gamma}\ \otimes\mathrm{id}_{\mathcal{K}})\widetilde{e}^{\,{\dagger}}\circ(-\widetilde{c}^{\,\prime}_{\gamma})\,\mathrm{mod}\,\widetilde{\mathcal{N}}^{\,{\dagger}}\langle 1\rangle\cap C_{i+1}(\widetilde{\mathcal{N}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
@@
We have also the following properties.